OpenStudy (anonymous):

WHICH OF THE FOLLOWING ARE EQUAL TO THE EXPRESSION BELOW ? CHECK ALL THAT APPLY 2*SQRT25 A.SQRT100 B.25 C.100 D.50 E.SQRT 50 F.10 ?? HELP PLZ ??

OpenStudy (anonymous):

Please, no caps lock...it's like screaming. \[2*(\sqrt{25})?\]

OpenStudy (anonymous):

\[2\sqrt{25}=2\times 5=10\]

OpenStudy (anonymous):

\[\sqrt{100}=10\]

OpenStudy (anonymous):

No What's\[\sqrt{25}\]

OpenStudy (anonymous):

and also \[10=10\]

OpenStudy (anonymous):

sqrt of 25 is 5

OpenStudy (anonymous):

whats 2 times 5?

OpenStudy (anonymous):

?? yes that is what i get. ten

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

F is also 10

OpenStudy (anonymous):

Sorry @satellite73 not tryin to steal your thunder, wasn't paying attention :)